OpenStudy (anonymous):

pentagon ABCDE ~ pentagon FGHIJ. Use the figures below to answer the following questions @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you're missing the figures

jimthompson5910 (jim_thompson5910):

or the following questions

OpenStudy (anonymous):

sorry i keep getting distracted

jimthompson5910 (jim_thompson5910):

that's ok

jimthompson5910 (jim_thompson5910):

AB corresponds to FG

jimthompson5910 (jim_thompson5910):

So FG/AB = 6/8 = 0.75 is the number you multiply each side length in the larger pentagon to get the corresponding side length in the smaller pentagon

OpenStudy (anonymous):

8?

jimthompson5910 (jim_thompson5910):

8 times 0.75 = 6

jimthompson5910 (jim_thompson5910):

oh wait, i read that wrong...one sec

jimthompson5910 (jim_thompson5910):

oh for some reason, it still works out...lucky coincidence but yes, you multiply by 0.75

OpenStudy (anonymous):

oh

jimthompson5910 (jim_thompson5910):

to go backwards, you divide by 0.75

jimthompson5910 (jim_thompson5910):

notice how FJ corresponds to AE

jimthompson5910 (jim_thompson5910):

and FJ = 6 So AE = FJ/0.75 = 6/0.75 = 8 which means AE = 8

OpenStudy (anonymous):

so then i had it right?

jimthompson5910 (jim_thompson5910):

yes, sry i didn't see that til now

OpenStudy (anonymous):

ok, cool and for #14 it would be 2.5?

OpenStudy (anonymous):

no no no it would be 3

jimthompson5910 (jim_thompson5910):

which side corresponds to IJ?

OpenStudy (anonymous):

CD, so it would make it 3? correct?

jimthompson5910 (jim_thompson5910):

this is why this problem is a bit confusing

jimthompson5910 (jim_thompson5910):

the second pentagon should be rotated slightly clockwise so that FG matches up with AB

jimthompson5910 (jim_thompson5910):

now tell me what ij corresponds to

OpenStudy (anonymous):

8

jimthompson5910 (jim_thompson5910):

did you do the rotation? this will make F the bottom most point

OpenStudy (anonymous):

yea

jimthompson5910 (jim_thompson5910):

so ij corresponds to ???

jimthompson5910 (jim_thompson5910):

which side from ABCDE

OpenStudy (anonymous):

de

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now multiply that length by 0.75 to find ij

OpenStudy (anonymous):

2.5

jimthompson5910 (jim_thompson5910):

try it again 5 times 0.75 = ???

jimthompson5910 (jim_thompson5910):

de = 5

OpenStudy (anonymous):

3.75

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

fg would be 4?

jimthompson5910 (jim_thompson5910):

fg corresponds to ???

OpenStudy (anonymous):

ab?

jimthompson5910 (jim_thompson5910):

good FG = 0.75 times AB

OpenStudy (anonymous):

6

jimthompson5910 (jim_thompson5910):

you nailed it

OpenStudy (anonymous):

thanks, this will be the last one i promise

jimthompson5910 (jim_thompson5910):

The two triangles are similar So AD/ED = BD/CD 35/16 = 25/x Now solve for x

OpenStudy (anonymous):

9.6?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

35/16 = 25/x 35x = 16*25

OpenStudy (anonymous):

7 6/7

jimthompson5910 (jim_thompson5910):

no it's not any of those answer choices

jimthompson5910 (jim_thompson5910):

x is the length of CD

jimthompson5910 (jim_thompson5910):

which isn't listed

OpenStudy (anonymous):

16*25=400

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

35x = 16*25 35x = 400

OpenStudy (anonymous):

dfsdfdddddddd

OpenStudy (anonymous):

sorry cat walked on the kep board... and i dont get wat i do next

OpenStudy (anonymous):

key*

jimthompson5910 (jim_thompson5910):

that's ok, divide both sides by 35

OpenStudy (anonymous):

45 divided by 35, and what else by 35?

OpenStudy (anonymous):

11.4

jimthompson5910 (jim_thompson5910):

ok now subtract this from 25

OpenStudy (anonymous):

13.6

jimthompson5910 (jim_thompson5910):

which is closest to which answer choice?

OpenStudy (anonymous):

13 4/7

OpenStudy (anonymous):

13.6 is closest to 13 6/7

jimthompson5910 (jim_thompson5910):

your first response is correct

jimthompson5910 (jim_thompson5910):

13.6 is closest to 13 4/7

jimthompson5910 (jim_thompson5910):

which is the answer

OpenStudy (anonymous):

13 4/7

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

thank you,... i was just wondering, but are u like a teacher or something?

jimthompson5910 (jim_thompson5910):

i'm a student learning to become one

OpenStudy (anonymous):

your are amazing

jimthompson5910 (jim_thompson5910):

why thank you, i appreciate the compliment

OpenStudy (anonymous):

your are verry verry welcome, you helped me out so much that i officaly have a B in geometry ( summerschool) i had failed it the first time idurring the year, i dont know what i would have done with out you...

jimthompson5910 (jim_thompson5910):

i'm glad i could help you out, you're very smart and it sounds like you're starting to get the hang of this stuff, which is great

OpenStudy (anonymous):

thank you :)

jimthompson5910 (jim_thompson5910):

you're welcome